An electric generator has e.m.f. of 240 V and an internal resistance of 1Ω if the current supplied by the generatoris 20A when the terminal voltage is 220 V, find the ratio of the power supplied to the power dissipated


1
Expert's answer
2021-06-21T11:38:36-0400

Let's first calculate power supplied to the generator:


"P_{supplied}=IE=20\ A\cdot240\ V=4800\ W."

Then, let's calculate the power dissipated by the generator:


"P_{dissipated}=IV=20\ A\cdot220\ V=4400\ W."

Finally, we can find the ratio of the power supplied to the power dissipated:


"\dfrac{P_{supplied}}{P_{dissipated}}=\dfrac{4800\ W}{4400\ W}=\dfrac{12}{11}."